# Form-From: A Design Space of Social Media Systems

### 08 Feb 2024 v1

@@ Abstract.
Social media systems are as varied as they are pervasive. They have been almost universally adopted for a broad range of purposes including work, entertainment, activism, and decision making. As a result, they have also diversified, with many distinct designs differing in content type, organization, delivery mechanism, access control, and many other dimensions. In this work, we aim to characterize and then distill a concise design space of social media systems that can help us understand similarities and differences, recognize potential consequences of design choice, and identify spaces for innovation. Our model, which we call Form-From, characterizes social media based on (1) the form of the content, either threaded or flat, and (2) from where or from whom one might receive content, ranging from spaces to networks to the commons. We derive Form-From inductively from a larger set of 62 dimensions organized into 10 categories. To demonstrate the utility of our model, we trace the history of social media systems as they traverse the Form-From space over time, and we identify common design patterns within cells of the model.
@@ 1. Introduction
We derived Form-From by first generating a larger set of 62 dimensions that aims to summarize the key types of differences between social media systems via an inductive process covering a large number of industry and research systems. We hierarchically grouped these 62 dimensions into ten categories: Discussion Structuring (simplified to “Form”), Delivery (simplified to “From”), Temporality, Audience, Membership, Moderation, Governance, Extensibility, Signaling, and Content. While the Form-From model captures what we argue are two key dimensions of variation across social media designs, these ten categories and 62 dimensions serve as a fuller treatment for comparing similar designs. For example, while Twitter, Mastodon, Threads, and Bluesky all occupy the threaded-network cell of the Form-From model, they differ in their approaches to Governance, Moderation, and Extensibility.
@@ 4.1.1. Dimension 1: The Form of Content
Our first dimension describes the way that social content is organized, or what form the content takes at a high level. The two possible Forms are threaded and flat. In a threaded system, the design emphasizes explicitly linking posts to each other, for example through replies. Example threaded designs include:
@@
  • Traditional forums, which organize replies chronologically around top-level posts and allow replying to individual posts.
@@
In some ways, the Form dimension correlates with the synchronous-asynchronous divide in the “time” dimension of the Johansen matrix, in that systems where people predominantly interact synchronously tend to appear in the Flat category, while systems where people predominantly interact asynchronously tend to appear in the Threaded category. However, we argue that in today’s social media, time is no longer a useful way to distinguish systems, as a vast number of systems are used today in a semi-synchronous fashion, making the synchronous-asynchronous divide more of a continuum. For instance, instant messaging, chat systems, and multi-user editors all appear in the synchronous side of the Johansen matrix (Johansen, 1988)—yet all of these systems can be used asynchronously and in many cases are actually predominantly used asynchronously today (Wang et al., 2017). Instead of focusing on usage, which can change as norms change, our focus on the fundamental shape of user-to-user interaction, as conveyed by the underlying design and data model, enables us to more cleanly separate systems and also to pinpoint what design choices facilitate certain kinds of interactions. For example, while both chat and email can be and are used asynchronously, the form of chat as a flat, chronological stream of distinct message units can flexibly suit more synchronous interaction (where replies are inferred by timing) than the form of email as a threaded, chronological system, which enables more asynchronous interaction (since one can reply to a long-ago message). [^8] Temporality does emerge as an orthogonal category in our 10-category expanded model, with many facets that go beyond synchronous-asynchronous interaction. Conversely, if any one channel is going to be used for multiple simultaneous conversations, some form of threading is generally required in order to keep those conversations from jumbling together.
@@
In this option, the design emphasizes explicitly linking posts to each other, for example through replies. Each piece of content can be a post sent by a user that kicks off a series of posts that are all linked back to the initial post. Meta’s system Threads, appropriately, is threaded by this definition.
There are different subspecies within this distinction. For example, threaded designs often have a form that might be called “headed”, as top-level content is treated differently from replies. Examples include Reddit and Facebook posts, where the top-level posts are a different beast than replies; for instance, only top-level posts can appear ranked in one’s feed. The “head” that kicks off discussion can also be its own distinct form of content, such as a news article, a document, or a video, that receives comments on it. Examples include YouTube and TikTok, where the initial content is a video, Instagram, where the initial content is an image, or comments sections of news organizations such as the New York Times, where the initial content is a news article. In general for these platforms, the only way to reach the comments hanging off a particular head is via that head content itself—e.g., the comments for a YouTube video can only be found on that video’s page. For these kinds of platforms, a further distinction involves whether there are one or more sub-threads that branch from individual replies. For instance, Reddit and Twitter allow for infinite branching of replies into sub-threads, while other platforms only allow a single thread or one or two levels of branching. [^9] This distinction surfaces as two of the possible options in the dimension “Threading/Partition of Replies” in our 62-dimension model (A.1.3).
@@
In contrast, flat designs do not connect different pieces of content through structured links. Instead, the design presents a stream of disconnected posts.
@@ 4.1.2. Dimension 2: Where Content Comes From
What determines which content you look at when you open the platform? Our second dimension is concerned with how content is delivered from one account to another. This includes users as both senders of content, where users have some ability in the system to designate how or to whom content is sent, as well as receivers of content, where users have some method of accessing content or determining how to locate a piece of content to which they have access.
@@
  • Spaces, where content is shared with a fixed group or locations, as in Slack, Microsoft Teams, or group chat;
  • Network, where content is shared based on friend or follow networks, as in Facebook, Twitter, and Instagram;
  • Commons, where content is shared with the intent that it be accessible to the entire platform, as in TikTok or Pinterest.
@@
One instructive example is group chat: while it may be tempting to consider group chat as a network-based design since the recipients are specific accounts, group chat is properly a space-based design because all members of the group chat “room” receive all messages—and many group chat platforms such as WhatsApp even allow or encourage naming the group chat like a channel and assign an admin to manage membership. Similarly, we would characterize directed mailing, messaging, and calling systems (e.g., email, SMS, Zoom) to be spaces.
@@
Traditional social network sites are the classic examples of social network delivery, as in the Facebook News Feed. Platforms that are centered around users subscribing to a creator’s or influencer’s content are also networks, such as Patreon or YouTube. Systems that implement “stories” such as Snapchat [^12] Though Snapchat’s Group chat surface would be considered a space. and Instagram Stories are also networks as they involve broadcasting short disappearing videos and images to one’s friends.
@@
Many platforms also blend commons designs with other delivery mechanisms. As of this writing, Twitter’s For You page is 50% centered around follows and 50% content from the commons based on a ranking algorithm; TikTok’s For You page is more skewed—principally a commons design but somewhat influenced by the follow network. Many social networks are beginning to augment their social network approach with the commons, claiming that this delivery mechanism increases engagement and content quality. For instance, Meta’s Threads app recently launched in mid-2023, with more of a commons-based delivery as it seeks to engage users who are still building a following graph.
@@ 4.1.3. Ideal Types for Each Cell in the 2x3 Model
  • Flat commons: a site for viewing algorithmically curated content, without the ability to comment or follow accounts
  • Threaded commons: a site for viewing and commenting on algorithmically curated content, without the ability to follow accounts
@@
The Form-From model captures two principal dimensions of variation across social media designs. However, a simple model cannot always capture dimensions of difference that matter to a particular system. For example, Mastodon, Bluesky, Threads, and Twitter all fall into the Threads + Network cell of the From-Form model, but each of them can also be described by important differences. For instance, Mastodon and Twitter differ strongly in their models of governance and extensibility, and all four differ strongly in their moderation strategy (Twitter’s is centralized and increasingly hands off, Threads’s is centralized and the most adherent to traditional content guidelines, Mastodon is decentralized to server administrators, and Bluesky places it in the hands of end users through middleware).
@@
To the mass media, the form of the content often defines a platform. Instagram was “the photo platform”, while Twitter was the “microblogging platform” and TikTok or Vine were the “short-form video platforms”. While we would argue that the content type is less determinative than these portrayals might imply, it still matters. What subset of content types are allowable on the platform: text, images, audio, videos, avatars, drawings, emojis? Platforms may also provide facilities for certain types of content, for example Instagram’s filters and TikTok’s stitch creation tool.
@@ 4.3. Full model: 62 dimensions
The full list of dimensions should be helpful as a checklist for designers: every social media design must make decisions on each dimension. Not making a decision means that the design does make a de-facto decision: to go with the typical or default value of the dimension. (For example, by default, platforms permanently keep archives.) However, defaulting can lead to mismatched expectations between users and platforms, and give rise to conflict later. So, we advocate that one practical use of our full 62-dimension model is as a design checklist for designers in the early stages of design. The full model is also useful for deep-dive inspections of closely related systems. For example, for research focused on social identity (Seering et al., 2018), comparative analyses across platform design decisions related to Membership and User and Group Signaling may be useful: What are the criteria for joining the platform or a group? Can a single person hold multiple accounts?
@@ 4.3.1. Dimension: Main Location of Discussion
A second option we observed for the dimension Main Location is rooms, where instead of one center where all discussion takes place, the space is first divided further into distinct addressable partitions, where users can enter the room to peruse content or post to it. Examples include Slack channels and Facebook Groups. One debate we had around this option is whether hashtags as popularized on Twitter, or tags in general, should count as “rooms”. Eventually we decided that while tags and rooms have some similar practical implications, they also have important differences, namely that while a piece of content may have many tags, they typically only have one main location (though can possibly be reposted or shared elsewhere). This discussion is also what led us to title this dimension “Main Location of Discussion”, recognizing that the same content can reside in different places. Instead, tags and hashtags appear in separate dimensions in this category such as Subscription (A.2.4) and Findability (A.2.6). A second debate we had was whether profile pages on systems such as Facebook and TikTok could be thought of as rooms, or whether these systems are centers. We eventually agreed that while one’s profile page has some properties in common with rooms, namely a unique address where content can be found, we still classify these systems as centers because their design is not primarily organized around navigating to profile pages in order to see content. However, a caveat is that early versions of some social networking systems like the Facebook Wall before the feed and their predecessors such as Blogger were room-like—one had to navigate to someone’s blog or wall to access any content.
@@
Next, we separated out cliques, which are like rooms, but instead of having an addressable name, cliques are distinguished by the membership of the shared space. An example of this distinction is found in Slack, where there can be channels with names like “#general”, but there are also multi-party direct messages, which have no name but are distinguished by the participants. An important difference between how rooms and cliques work is shared history—as users get added or removed from a channel, the historical content stays intact, much like one would expect of objects in a physical room where people can walk in and out. Indeed, a room still exists even if it is entirely empty of people. In a clique, however, if one adds a new user to a multi-party message, it is now a new clique with its own distinct history. We have found this subtle distinction leads to some interesting and unexpected behavior in different systems. For instance, we found through investigation that Facebook Messenger’s multi-party messages are closer to rooms than cliques, e.g., members of a multi-party message can add new people to the chat who can see history from before they were there, and members can also rename the shared chat as well as appoint admins. From our investigation, almost all group chat systems actually are designed more like rooms.
@@
The final option is mailboxes, where there is no shared location for discussion but all content gets delivered to personal inboxes. Email is the canonical example, where even if an email is removed from one participant’s inbox, it is still available on all other participants’ inboxes, as they each have their own copy. In addition, each user can choose to customize and use their desired email client as opposed to a unified experience, which would be more akin to traveling to a shared center or room. We discussed among ourselves how technically, most email systems are now hosted on a centralized service via the cloud as opposed to physically on a client device. Nevertheless the design of some email systems is still reminiscent of physical mail deliveries—but there are increasingly systems that would traditionally act like mailboxes, such as direct messaging, that now have room features in their designs.
@@
Our iterative discussions helped us isolate the important dimensions that made their way to the 2x3 model. In the case of this dimension, we determined that the design distinctions today between rooms, cliques, and mailboxes are overall subtle, so when we went to summarize our dimensions in our higher level 2x3 model, all of the systems with these values actually end up in the Spaces option of the “From” dimension. In contrast, systems that have one single center end up not having much of a spatial metaphor, due to the primacy of algorithms curating a central feed reducing the need for much navigation. Instead, these systems became spread out between the options of Network and Commons according to their placement on lower level dimensions such as Ranking of Posts (A.2.2), or in what order do posts appear, combined with Maximal Audience (A.4.2), or to what posts do users have access.

### 24 Mar 2024 v3

@@ Abstract.
Social media systems are as varied as they are pervasive. They have been almost universally adopted for a broad range of purposes including work, entertainment, activism, and decision making. As a result, they have also diversified, with many distinct designs differing in content type, organization, delivery mechanism, access control, and many other dimensions. In this work, we aim to characterize and then distill a concise design space of social media systems that can help us understand similarities and differences, recognize potential consequences of design choices, and identify spaces for innovation. Our model, which we call Form-From, characterizes social media based on (1) the form of the content, either threaded or flat, and (2) from where or from whom one might receive content, ranging from spaces to networks to the commons. We derive Form-From inductively from a larger set of 62 dimensions organized into 10 categories. To demonstrate the utility of our model, we trace the history of social media systems as they traverse the Form-From space over time, and we identify common design patterns within cells of the model.
@@ 1. Introduction
We derived Form-From by first generating a larger set of 62 dimensions that aims to summarize the key types of differences between social media systems via an inductive process covering a large number of industry and research systems. We hierarchically grouped these 62 dimensions into ten categories: Discussion Structuring (simplified to “Form”), Delivery (simplified to “From”), Temporality, Audience, Membership, Moderation, Governance, Extensibility, Signaling, and Content. While the Form-From model captures what we argue are two key dimensions of variation across social media designs, these ten categories and 62 dimensions serve as a fuller treatment for comparing similar designs. For example, while Twitter, Mastodon, and Bluesky all occupy the threaded-network cell of the Form-From model, they differ in their approaches to Governance, Moderation, and Extensibility.
@@ 4.1.1. Dimension 1: The Form of Content
Our first dimension describes the way that social content is organized, or what form the content takes at a high level. The two Forms are threaded and flat. In a threaded system, the design emphasizes explicitly linking posts to each other, for example through replies. Example threaded designs include:
@@
  • Traditional forums, which organize top-level posts chronologically and allow replies to them.
@@
In some ways, the Form dimension correlates with the synchronous-asynchronous divide in the “time” dimension of the Johansen matrix, in that systems where people predominantly interact synchronously tend to appear in the Flat category, while systems where people predominantly interact asynchronously tend to appear in the Threaded category. However, we argue that in today’s social media, time has become a less useful way to distinguish systems, as a vast number of systems are used today in a semi-synchronous fashion, making the synchronous-asynchronous divide more of a continuum. For instance, instant messaging, chat systems, and multi-user editors all appear in the synchronous side of the Johansen matrix (Johansen, 1988)—yet all of these systems can be used asynchronously and in many cases are predominantly used asynchronously today (Wang et al., 2017). Instead of focusing on usage, which can change as norms change, our focus on the fundamental shape of user-to-user interaction, as conveyed by the underlying design and data model, enables us to more cleanly separate systems and also to pinpoint what design choices facilitate certain kinds of interactions. For example, while both chat and email can be and are used both synchronously and asynchronously, they seem optimal for opposite modalities. Chat works well in a synchronous setting: since each post is typically replying (if at all) to the one immediately previous, there is little need for presenting complex reply threads and a flat chronological stream offers a simpler user experience. In contrast, email works well for asynchronous conversations, because the explicitly represented threading structure supports navigation of multiple distinct conversations overlapping in time. [^8] Temporality does emerge as an orthogonal category in our 10-category expanded model, with many facets that go beyond synchronous-asynchronous interaction. Conversely, if any one channel is going to be used for multiple simultaneous conversations, some form of threading is generally required in order to keep those conversations from jumbling together.
@@
In this option, the design emphasizes explicitly linking posts to each other, for example through replies. Each piece of content can be a post sent by a user that kicks off a series of posts that are all linked back (possibly indirectly) to the initial post. Meta’s system Threads, appropriately, is threaded by this definition.
There are different subspecies within this distinction. For example, threaded designs often have a form that we call headed, as top-level content is treated differently from replies. Examples include Reddit and Facebook posts, where the top-level posts are quite distinct from replies; for instance, only top-level posts can appear ranked in one’s feed. The “head” that kicks off discussion can also be its own distinct form of content, such as a news article, a document, or a video, that receives comments on it. Examples include YouTube and TikTok, where the initial content is a video, Instagram, where the initial content is an image, or comments sections of news organizations such as the New York Times, where the initial content is a news article. In general for these platforms, the only way to reach the comments hanging off a particular head is via that head content itself—e.g., the comments for a YouTube video can only be found on that video’s page. For these kinds of platforms, a further distinction involves whether there are one or more sub-threads that branch from individual replies. For instance, Reddit and Twitter allow for infinite branching of replies into sub-threads, while other platforms only allow a single thread or one or two levels of branching. [^9] This distinction surfaces as two of the possible options in the dimension “Threading/Partition of Replies” in our 62-dimension model (A.1.3).
@@
In contrast, flat designs do not connect different pieces of content through structured links. Instead, the design presents a stream of unconnected posts.
@@ 4.1.2. Dimension 2: Where Content Comes From
What determines which content you see when you visit the platform? Our second dimension is concerned with how content is delivered from one account to another. This includes users as both senders of content, where users have some ability in the system to designate how or to whom content is sent, as well as receivers of content, where users have some method of accessing content or determining how to locate a piece of content to which they have access.
@@
  • Spaces, where a user receives content that has been submitted to certain virtual locations, channels, or groups that the user “joins”, as in Slack, Microsoft Teams, or group chat;
  • Network, where one receives content from specified friends or follow networks, as in Facebook, Twitter, and Instagram;
  • Commons, where the user notionally draws their content from the entire platform, as in TikTok or Pinterest.
@@
One instructive example is group chat: while it may be tempting to consider group chat as a network-based design since the recipients are specific accounts, group chat is properly a space-based design because all members of the group chat “room” receive all messages—and many group chat platforms such as WhatsApp even allow or encourage naming the group chat like a channel and assign an admin to manage membership. Similarly, we would characterize directed mailing, messaging, and calling systems (e.g., email, SMS, Zoom) to be spaces. For example, although email delivery is typically individual to individual, suggesting a network model, when a user composes a message to a group of recipient, followup conversation typically makes use of “Reply All” to engage all the original recipients. So the original sender essentially creates a space in which all follow-up messages are typically delivered.
@@
Traditional social network sites are the classic examples of social network delivery, as in the Facebook News Feed. Platforms that are centered around users subscribing to a creator’s or influencer’s content are also networks, such as Patreon, Substack, or YouTube. Systems that implement “stories” such as Snapchat [^12] Though Snapchat’s Group chat surface would be considered a space. and Instagram Stories are also networks as they involve broadcasting short disappearing videos and images to one’s friends.
@@
Many platforms also blend commons designs with other delivery mechanisms. As of this writing, Twitter’s For You page is 50% centered around follows and 50% content from the commons based on a ranking algorithm. TikTok’s For You page is more skewed—principally a commons design but somewhat influenced by the follow network. Many social networks are beginning to augment their social network approach with the commons, claiming that this delivery mechanism increases engagement and content quality. For instance, Meta’s Threads app recently launched in mid-2023, with more of a commons-based delivery as it seeks to engage users who are still building a following graph.
@@ 4.1.3. Ideal Types for Each Cell in the 2x3 Model
  • Flat commons: a site for viewing algorithmically curated content, without the ability to reply or follow accounts
  • Threaded commons: a site for viewing and replying to algorithmically curated content, without the ability to follow accounts
@@
The Form-From model captures two principal dimensions of variation across social media designs. However, a simple model cannot always capture dimensions of difference that matter to a particular system. For example, Mastodon, Bluesky, and Twitter all fall into the Threads + Network cell of the From-Form model, but each of them can also be described by important differences. For instance, Mastodon and Twitter differ strongly in their models of governance and extensibility, and all three differ strongly in their moderation strategy (Twitter’s is centralized and increasingly hands off, Mastodon is decentralized to server administrators, and Bluesky places it in the hands of end users through middleware).
@@
To the mass media, the type of content often defines a platform. Instagram was “the photo platform”, while Twitter was the “microblogging platform” and TikTok or Vine were the “short-form video platforms”. While we would argue that the content type is less determinative than these portrayals might imply, it still matters. What subset of content types are allowable on the platform: text, images, audio, videos, avatars, drawings, emojis? Platforms may also provide facilities for certain types of content, for example Instagram’s filters and TikTok’s stitch creation tool.
@@ 4.3. Full model: 62 dimensions
The full list of dimensions should be helpful as a checklist for designers: every social media design must make decisions on each dimension. Not making a decision means that the design does make a de-facto decision to go with the typical or default value of the dimension. (For example, by default, platforms permanently keep archives.) However, defaulting can lead to mismatched expectations between users and platforms, and give rise to conflict later. So, we advocate that one practical use of our full 62-dimension model is as a design checklist for designers in the early stages of design. The full model is also useful for deep-dive inspections of closely related systems. For example, for research focused on social identity (Seering et al., 2018), comparative analyses across platform design decisions related to Membership and User and Group Signaling may be useful: What are the criteria for joining the platform or a group? Can a single person hold multiple accounts?
@@ 4.3.1. Dimension: Main Location of Discussion
A second option we observed for the dimension Main Location is rooms—a special case of spaces in the From dimension—where instead of one center where all discussion takes place, the space is first divided further into distinct addressable partitions, where users can enter the room to peruse content or post to it. Examples include Slack channels and Facebook Groups. One debate we had around this option is whether hashtags as popularized on Twitter, or tags in general, should count as “rooms”. Eventually we decided that while tags and rooms have some similar practical implications, they also have important differences, namely that while a piece of content may have many tags, they typically only have one main location (though can possibly be reposted or shared elsewhere). This discussion is also what led us to title this dimension “Main Location of Discussion”, recognizing that the same content can reside in different places. Instead, tags and hashtags appear in separate dimensions in this category such as Subscription (A.2.4) and Findability (A.2.6). A second debate we had was whether profile pages on systems such as Facebook and TikTok could be thought of as rooms, or whether these systems are centers. We eventually agreed that while one’s profile page has some properties in common with rooms, namely a unique address where content can be found, we still classify these systems as center-based because their design is not primarily organized around navigating to profile pages in order to see content. However, a caveat is that early versions of some social networking systems like the Facebook Wall before the feed and their predecessors such as Blogger were room-like—one had to navigate to someone’s blog or wall to access any content.
@@
Next, we separated out cliques, which are like rooms, but instead of having an addressable name, cliques are distinguished by the membership of the shared space. An example of this distinction is found in Slack, where there can be channels with names like “#general”, but there are also multi-party direct messages, which have no name but are distinguished by the participants. An important difference between how rooms and cliques work is shared history—as users get added or removed from a channel, the historical content stays intact, much like one would expect of objects in a physical room where people can walk in and out. Indeed, a room still exists even if it is entirely empty of people. In a clique, however, if one adds a new user to a multi-party message, it is now a new clique with its own distinct history. We have found this subtle distinction leads to some interesting and unexpected behavior in different systems. For instance, we found through investigation that Facebook Messenger’s multi-party messages are closer to rooms than cliques, e.g., members of a multi-party message can add new people to the chat who can see history from before they were there, and members can also rename the shared chat as well as appoint admins. Facebook’s user experience confuses this issue: if a user composes a message to the group by listing their names, and they have messaged the exact same group previously, then Facebook drops them into the preexisting group channel where previous messages can be seen. This creates the illusion that communication is based on the clique. However, if the group’s old room gets renamed, then messaging the same group results in the creation of a new, different channel for the new conversation. From our investigation, almost all group chat systems actually are designed more like rooms.
@@
The final option is mailboxes, where there is no shared location for discussion but all content gets delivered to personal inboxes. Email is the canonical example, where even if an email is removed from one participant’s inbox, it is still available on all other participants’ inboxes, as they each have their own copy. In addition, each user can choose to customize and use their desired email client as opposed to the shared experience of a common center or room. We discussed among ourselves how technically, most email systems are now hosted on a centralized service via the cloud as opposed to physically on a client device. Nevertheless the design of some email systems is still reminiscent of physical mail deliveries. There are also increasingly systems that would traditionally act like mailboxes, such as direct messaging, that now have room features in their designs.
@@
Our iterative discussions helped us isolate the important dimensions that made their way to the 2x3 model. In the case of this dimension, we determined that the design distinctions today between rooms, cliques, and mailboxes are overall subtle, so when we went to summarize our dimensions in our higher level 2x3 model, all of the systems with these values actually end up in the Spaces option of the “From” dimension. In contrast, systems that have one single center end up not having much of a spatial metaphor, due to the primacy of algorithms curating a central feed reducing the need for much navigation. Instead, these systems became spread out between the options of Network and Commons according to their placement on lower level dimensions such as Ranking of Posts (A.2.2), or in what order posts appear, combined with Maximal Audience (A.4.2), or to what posts do users have access.